OpenStudy (mony01):

Find the value(s) of x between pi/2 and 3pi/2 for which F(x) has a global maximum and global minimum.

OpenStudy (mony01):

\[F(x)\int\limits_{\pi/2}^{x}\frac{ \sin(t) }{ t}dt\]